OpenStudy (anonymous):

Find (f^-1)'(a) of: 2x^3+3x^2+7x+4

OpenStudy (anonymous):

Can you do the derivative of f(x) or do the inverse first and figure the Derivative?

OpenStudy (anonymous):

\[f'(x) = 6x ^{2}+6x+7\]

OpenStudy (anonymous):

Would it be easier ("nicer") to do the inverse of the f'(x) or just the inverse of f(x)? I think it will be ugly either way?

zepdrix (zepdrix):

Hmm I actually haven't seen this type of problem before :D lol Lemme think about it a few minutes... My initial guess is that we probably have to take the inverse first, get it back in terms of x (somehow..) THEN differentiate.

OpenStudy (anonymous):

Ok, I will keep trying it also. I have done just that: \[y=2x ^{3}+3x ^{2}+7x+4\] \[x=2y ^{3}+3y ^{2}+7y+4\] Then solve. That is what i have been doing.

OpenStudy (anonymous):

Also tried the same above but with the derivative.

OpenStudy (anonymous):

My chapter is eventually working into Exponential,Logarithmic, and Inverse Trig Functions.

OpenStudy (anonymous):

\[x=2y ^{3}+3y ^{2}+7y+4\] \[x-4 = y(2y ^{2}+3y+7)\] \[x-4=y(2y+1)(y+3)\]

OpenStudy (anonymous):

?

zepdrix (zepdrix):

\[\large \left[f^{-1}\right]'(a)=\frac{1}{f'(f^{-1}(a))}\]I keep coming across this equation.. still trying to make sense of it though :D heh

OpenStudy (anonymous):

DOH SORRY, totally forgot. a=4

zepdrix (zepdrix):

Oh hmm maybe we can do something with that :)

OpenStudy (anonymous):

Which you still need to find the inverse though ... right?

zepdrix (zepdrix):

seems like it XD lol that seems to be the tricky part, hmmm

OpenStudy (anonymous):

Of course it will be tricky :-) Calculus = 1% Calc and 99% Algebra. But that 1% Calculus make the Algebra "tricky".

OpenStudy (anonymous):

Is it "legal" to do the following? \[x=2y ^{3}+3y ^{2}+7y+4\] \[x ^{1/3}=2(\sqrt[3]{3y ^{2}+7y+4}\]

zepdrix (zepdrix):

Where did the y^3 go? You were attempting to take the cube root of that separately from the rest of the thing? Mmm yah that doesn't quite work the way you want it to :C

OpenStudy (anonymous):

figured :-)

OpenStudy (anonymous):

How about what I did up above, would that get us anywhere?

zepdrix (zepdrix):

Maybe yes, let's think about this a second...\[\large x=2y^3+3y^2+7y+4\]Here is the inverse function of f, written in terms of Y. If we were ABLE to write it in terms of x, we would be plugging in a=4 for the X value right? Let's try plugging in 4 for the X value right now, maybe that will get us somewhere.\[\large 4=2y^3+3y^2+7y+4 \qquad \rightarrow \qquad 0=2y^3+3y^2+7y\]

OpenStudy (anonymous):

hmmmmmm \[f ^{-1}(y) = x <---> f(x)=y\] Could we put a into f(x) and that would give us y?

zepdrix (zepdrix):

Hmm that's an interesting idea :o

OpenStudy (anonymous):

\[f(4)=208\] \[f ^{-1}(208)=4\] \[(f ^{-1})'(a)=\frac{ 1 }{ f'(f ^{-1}(4)) }\] \[(f ^{-1})'(a) = \frac{ 1 }{ f'(208) }\] \[(f ^{-1})'(a)=\frac{ 1 }{ 260839 }\] But .... Not even close t the answer in the back.

zepdrix (zepdrix):

What's the back say? :o

OpenStudy (anonymous):

Heh \[\frac{ 1 }{ 7 }\]

OpenStudy (anonymous):

Close in some aspect :-)

zepdrix (zepdrix):

lol :3

OpenStudy (anonymous):

I typed this into Wolfram: inverse of 2x^3+3x^2+7x+4 And got this:

OpenStudy (anonymous):

I hope that this is not what I need to do. That is ....... not pretty.

zepdrix (zepdrix):

No I'm sure it's not what you need to do. There is clearly some little trick that we're just not seeing. :\

OpenStudy (anonymous):

And everyone who pops on to take a look runs screaming.

zepdrix (zepdrix):

@Zarkon is a pretty smart dude :d Maybe he'll take a look at it if he has time. :D I might have to give up on this one :C grr

OpenStudy (anonymous):

That's ok. I appreciate all the help. I may try to send a message and see what happens. Thank you.

zepdrix (zepdrix):

Well I did the little @ symbol thing. That is the best way to try and get someones attention. :D It's like paging someone. It lets them know that you called, and they can click on the page to come directly to your thread. So it's a handy little tool to use :D

OpenStudy (anonymous):

Didn't know that. Thanks for the little nugget of info.

OpenStudy (anonymous):

Hello

OpenStudy (anonymous):

@jim_thompson5910 Help?

OpenStudy (anonymous):

@helder_edwin Help? Inverse Derivatives

OpenStudy (anonymous):

Hello

OpenStudy (anonymous):

Thank you for coming and taking a look.

jimthompson5910 (jim_thompson5910):

a = 4 correct?

OpenStudy (anonymous):

Correct.

jimthompson5910 (jim_thompson5910):

when you plug in x = 0, you get f(x) = 4 so f^-1(4) = 0, agreed?

OpenStudy (anonymous):

agreed .....

jimthompson5910 (jim_thompson5910):

so... [ f^-1]' (a) = 1/[ f' ( f^-1(a) ) ] [ f^-1]' (4) = 1/[ f' ( f^-1(4) ) ] [ f^-1]' (4) = 1/[ f' ( 0 ) ] [ f^-1]' (4) = 1/7

OpenStudy (anonymous):

so that would make: \[\frac{ 1 }{ f'(0) }\]

jimthompson5910 (jim_thompson5910):

f(x) = 2x^3+3x^2+7x+4 f ' (x) = 6x^2+6x+7 f ' (0) = 6(0)^2+6(0)+7 f ' (0) = 7

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

yes. So completly wrong plugging 4 in and solving the equation.

jimthompson5910 (jim_thompson5910):

what do you mean?

OpenStudy (anonymous):

Posted this earlier: hmmmmmm f−1(y)=x<−−−>f(x)=y Could we put a into f(x) and that would give us y?

jimthompson5910 (jim_thompson5910):

it's the other way around, you plug in y = a and solve for x

jimthompson5910 (jim_thompson5910):

to find f^-1(a)

OpenStudy (anonymous):

So I did f(x) = 2(4)^3+3(4)^2+7(4)+4 When I should of done: 4=.............

jimthompson5910 (jim_thompson5910):

if you plugged in x = 4 into f(x), you'd get 208, but that's way off of what we want

jimthompson5910 (jim_thompson5910):

yeah exactly

jimthompson5910 (jim_thompson5910):

luckily 4 is the y-intercept, so it's kinda easy to solve for (if you know what to look for)

OpenStudy (anonymous):

ohhh man way to easy, thank you form coming to help.

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

Thank you for coming to take a look. Made it way to complecated.